OpenStudy (anonymous):

How do you solve this? 4x - 5y = 20

OpenStudy (anonymous):

@Choe_Caday Do you want to know the numbers to fill in for the variables?

OpenStudy (anonymous):

@lizarcNeris yes please ☺

OpenStudy (anonymous):

Okay my answer is x=15 and y=8 4*15=60 5*8=40~60-40=20

OpenStudy (anonymous):

Okay! TYSM! ♥

OpenStudy (anonymous):

No problem ^^